40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  PSL  Not logged in ELOG logo
Message ID: 1510     Entry time: Tue Sep 9 23:22:41 2014
Author: Tara, Evan 
Type: DailyProgress 
Category: FSS 
Subject: South 200 kHz oscillation; cavity pole 

We locked the south cavity using the north TTFSS. 200 kHz oscillation is still present, so whatever this is probably doesn't reside in the TTFSS box.

Tara and I also took cavity pole measurements using the EOAM and two PDA10CSs, one placed before the cavity (monitoring the light rejected from the post-EOAM PBS) and one after the cavity (on the ISS breadboard). The HP4395A was used to drive the EOAM, and we then took the transfer function which takes the pre-cavity PD voltage to the transmission PD voltage. I will fit these later, but the poles appear to be consistent with the values tabulated in ctn:1475.

Attachment 1: algaasFinesse.pdf  84 kB  Uploaded Wed Sep 10 00:56:38 2014  | Hide | Hide all
algaasFinesse.pdf
Attachment 2: algaasFinesse.zip  8 kB  Uploaded Wed Sep 10 11:45:16 2014
ELOG V3.1.3-